# Mitigating Message Imbalance in Fraud Detection with Dual-View Graph Representation Learning

### v1

@@ 3.1 Definition
Fraud detection on multi-relation graph Given a graph $G=\{\mathcal{V},\mathbf{X},\{{{\mathcal{E}}_{r}}\}|_{r=1}^{R},\mathcal{Y}\}$,which ${{v}_{i}}\in\mathcal{V}$ denotes the node, $\mathbf{X}$ denotes the features of the node, $\mathcal{R}$ denotes the kind of relationship, ${{e}_{i,j,r}}\in\mathcal{E}$ denotes the two nodes ${{v}_{i}}$ and ${{v}_{j}}$ are connected under the relationship ${r}$. Graph-based fraud detection is a semi-supervised binary classification task. Only a small number of nodes are labeled (denoted by $\mathcal{Y}$) on the graph $G$. Each node has a binary label. ${{y}_{i}}=0$ denotes a benign user, ${{y}_{i}}=1$ denotes a fraudster.
@@ 4.1 Topological Message Reachability
where $c$ represents the labeled categories 00 and $1$ of the labeled nodes, $\alpha^{r}$ denotes the probability that class $c$ is restarted by a randomized wandering under a relational subgraph; $A^{{}^{\prime}}_{r}=A_{r}D^{-1}$, $A$ denotes the adjacency matrix of the subgraph of relation $r$ and $D$ is the degree matrix. Additionally, $I^{r}_{c}\in R^{n}$ is the transmission vector, defined as follows:
$I^{r}_{c}=\begin{cases}\frac{1}{\left|\mathcal{V}_{L}^{c}\right|},&\text{ if }y_{i}=c\\ $0$,&otherwise\end{cases}$ (2)
@@ 4.2 Local Confounding Debiasing
$\mathcal{L}_{LCD}=$ $\sum_{j=1}^{p}\left(\gamma^{\mathrm{T}}\cdot\mathbf{h}_{i,m}^{\mathrm{T}}\left|\Lambda_{\mathbf{w}}\mathbf{h}_{i,-m}/n\right.\right.$ (12)
@@
[…]

### v2

@@ 3.1 Definition
Fraud detection on multi-relation graph Given a graph $G=\{\mathcal{V},\mathbf{X},\{{{\mathcal{E}}_{r}}\}|_{r=1}^{R},\mathcal{Y}\}$,which ${{v}_{i}}\in\mathcal{V}$ denotes the node, $\mathbf{X}$ denotes the features of the node, $\mathcal{R}$ denotes the kind of relationship, ${e}_{i,j,r}$$\in\mathcal{E}$ denotes the two nodes ${v}_{i}$} and ${v}_{j}$} are connected under the relationship r{r}. Graph-based fraud detection is a semi-supervised binary classification task. Only a small number of nodes are labeled (denoted by $\mathcal{Y}$) on the graph $G$. Each node has a binary label. ${{y}_{i}}=0$ denotes a benign user, ${{y}_{i}}=1$ denotes a fraudster.
@@ 4.1 Topological Message Reachability
where $c$ represents the labeled categories 0 and $1$ of the labeled nodes, $\alpha^{r}$ denotes the probability that class $c$ is restarted by a randomized wandering under a relational subgraph; $^{{}^{\prime}}_{r}=A_{r}D^{-1}, AA$ denotes the adjacency matrix of the subgraph of relation $r$ and $D$ is the degree matrix. Additionally, $I^{r}_{c}\in R^{n}$ is the transmission vector, defined as follows:
[…]
@@ 4.2 Local Confounding Debiasing
[…]
@@
$\text{ s.t. }\mathbf{w}\geq 0$
where $\mathbf{h}_{i,m}$ denotes the node representation $m$ of the variables in $\mathbf{h}_{i}$, and $\mathbf{h}_{i,-m}$ is the node representation $\mathbf{h}_{i}$ of all the variables in $\mathbf{h}_{i}$ except $m$. The first term represents the computation of the differences between different variables in the node representation and applies weights $\gamma$ to these variables to amplify their association with the true identity of the node. The second term indicates stability by enforcing regularization. The third serves to prevent the sample weights from going to zero. The  $\text{ s.t. }\mathbf{w}\geq 0$ is a non-negative constraint. To reduce computational complexity, this regularization is used at the last layer of the linear transformation.
@@ Acknowledgments
Appendix A Experiments
A.1 Dataset
The statistics of the dataset are shown in table 3. In addition, a description of the dataset is as follows:
  • YelpChi: collects reviews (spam) and recommendations (legitimate) on Yelp platforms about hotels and restaurants, with the following three main relationships: 1) R-U-R: reviews posted by the same user are connected, 2) R-T-R: reviews posted under the same month based on the same product are connected, and 3) R-S-R: reviews under the same product with the same star rating are connected.
[…]
  • Comp: records the financial statements of Chinese companies. It contains three relationships: 1) C-I-C: connects companies with investment relationships, 2) C-P-C: connects companies and their disclosed customers, and 3) C-S-C: connects companies and their disclosed suppliers.
Dataset #Nodes Fraud(%) Relation #Edges
R-U-R 98,630
YelpChi 45,954 14.5% R-T-R 1,147,232
R-S-R 6,805,486
U-P-U 351,216
Amazon 11,944 9.5% U-S-U 7,132,958
U-S-U 2,073,474
C-I-C 5,686
Comp 5,317 10.5% C-S-C 760
C-P-C 1,043
Table 3: The Statistics of Datasets. # represents quantity.
A.2 Baseline
A description of the backbone model used is as follows:
  • GCN Kipf and Welling [2017]: a graph convolutional network that generates final node embeddings by aggregating information from first-order neighbors.
  • GAT Velickovic et al. [2017]: a graph attention network with neighbor weights computed by the attention mechanism and eventually aggregated based on the weights.
  • GraphSAGE Hamilton et al. [2017]: an inductive framework that samples and aggregates neighboring nodes to generate node representations.
  • CARE-GNN Dou et al. [2020]: a camouflage-resistant graph neural network for neighbor selection using label similarity metric and reinforcement learning.
  • PC-GNN Liu et al. [2021]: a fraud detection method that balances node distribution using labels and neighborhood samplers.
  • FRAUDER Zhang et al. [2021]: a fraud detection effort that aggregates the differences between neighbors and handles class imbalances.
  • GAGA Wang et al. [2023]: a fraud detection method based on a Transformer capturing semantic information to deal with the low congruence problem.
  • COFD Hu et al. [2023]: a fraud detector to detect collaborative fraud by focusing on the comparison of first and second order node pair structures.
  • ConsisGAD Chen et al. [2024]: a fraud detector that mitigates class imbalance by performing domain consistency augmentation for homophilic distribution of normal and fraudulent nodes.
A.3 Matrics
The meanings of the indicators we use are as follows:
  • Recall indicates the proportion of correctly predicted positive instances among all actual positive instances in the sample; a higher value implies better performance.
  • AUC provides a reliable evaluation of the classifier’s overall performance, particularly effective in patterns with imbalanced samples.
  • F1-score represents the harmonic mean of precision and recall, with a higher value indicating increased precision and recall while minimizing variance.
A.4 Additional results
T-Finance [^1] Jianheng Tang, et al. Rethinking Graph Neural Networks for Anomaly Detection. ICML’22. is a single-relationship fraud graph designed to identify anomalous accounts within a trading network. To assess the method’s utility on a larger dataset, we compared it against selected baselines. As shown in  4, MimbFD consistently outperforms the baselines, demonstrating its ability on the larger dataset effectively mitigate the imbalanced propagation of supervisory messages in fraud detection scenarios.
Model AUC Recall F1
GCN 88.33±1.06 51.24±9.87 61.37±3.71
ReNode 42.79±28.85 79.40±5.84 46.5±8.2
CARE-GNN 91.76±0.32 78.11±0.36 73.54±0.65
GAGA 95.53±0.45 87.78±0.58 90.65±1.25
PC-GNN - - -
ConsisGAD 96.61±0.07 78.32±1.59 91.51±0.05
MimbFD 96.69±0.44 88.67±0.39 92.05±0.33
Table 4: Performance Comparison on T-Finance. Bold: the best of baselines.